# Explaining Grokking and Information Bottleneck through Neural Collapse Emergence

### v1

@@ 1 Introduction
Deep neural networks (DNNs) have demonstrated remarkable success across a variety of tasks, including computer vision, natural language processing, and reinforcement learning, yet their training dynamics under gradient descent often reveal unexpected behavior. In particular, when training continues beyond the point where the training loss has been sufficiently reduced, several intriguing late-phase phenomena have been reported. One such phenomenon is grokking (Power et al., 2022), where models initially converge to an overfitting solution that perfectly fits the training data but fails to generalize to unseen data. However, when training continues for a sufficiently long time, the models unexpectedly generalize. Another example is the information bottleneck (IB) framework (Tishby et al., 2000; Tishby & Zaslavsky, 2015; Shwartz-Ziv & Tishby, 2017), an information-theoretic perspective on representation learning in DNNs that formalizes the goal of retaining task-relevant information while compressing the task-irrelevant input information. One particularly intriguing observation here is that DNNs do not move directly toward an IB-optimal solution; instead, they first enter a fitting phase where they memorize the training data, followed by a later compression phase in the later training stage during which task-irrelevant input information is discarded.
@@
We consider a $K$-class classification problem with a dataset $S=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ consisting of $N$ examples, where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ is the input and $y_{i}\in[K](\coloneqq\{1,2,\dots,K\})$ is the class label. The input domain is ${\mathcal{X}}\subseteq\mathbb{R}^{d}$, and ${\bm{y}}_{i}$ denotes the one-hot encoding label in $\{0,1\}^{K}$. Let ${\bm{X}}\in\mathbb{R}^{d\times N}$ and ${\bm{Y}}\in\{0,1\}^{K\times N}$ denote the training inputs and one-hot labels, respectively. We use $S_{c}=\{({\bm{x}}_{i},y_{i})\in S\mid y_{i}=c\}$ to denote the subset of class-$c$ examples, and, by abuse of notation, also the corresponding index set $\{i\in[N]\mid y_{i}=c\}$. We denote its cardinality as $n_{c}=|S_{c}|$. Let $X$ and $Y$ denote the random variables representing the input and label, respectively. We write $p_{X,Y}$ for the joint distribution over $(X,Y)$, and $p_{X}$ and $p_{Y}$ for their marginal distributions. We simply write $p$ when the variables are clear from context. Let $I(X;Y)$ be the mutual information (MI) between random variables $X$ and $Y$. We denote the multivariate Gaussian distribution with mean $\bm{\mu}$} and covariance matrix $\bm{\Sigma}$} as $N({\bm{\mu}},{\bm{\Sigma}})$. We use $\log(\cdot)$ to denote the natural logarithm. For linear-algebraic notation, for a matrix $\bm{A}$}, we use $\|{\bm{A}}\|_{2}$ to denote the spectral norm, and $\|{\bm{A}}\|_{2,1}$ to denote the $(2,1)$ matrix norm, defined as $\|{\bm{A}}\|_{2,1}=\sum_{i}\|{\bm{a}}_{i}\|_{2}$, where ${\bm{a}}_{i}$ is the $i$-th column of $\bm{A}$}. Finally, we use $\langle{\bm{a}},{\bm{b}}\rangle={\bm{a}}^{\top}{\bm{b}}$ to denote the standard Euclidean inner product.
@@ 2 Related Work
On the empirical side, several studies attempted to explain the cause of grokking in terms of the parameter compression (Liu et al., 2023a; Varma et al., 2023) and some complexity measures (Nanda et al., 2023; Liu et al., 2023b; Humayun et al., 2024; DeMoss et al., 2025). Other studies sought to relate grokking to other training-related concepts, such as double descent (Davies et al., 2023; Huang et al., 2024) and optimization stability (Thilak et al., 2022). On the theoretical side, a high-dimensional limit of the linear model was analyzed in the setting of regression (Levi et al., 2024) and binary classification (Beck et al., 2025). There are also studies analyzing two-layer networks with XOR data (Xu et al., 2024) and mean-field analysis (Rubin et al., 2024). The most closely related line of work focuses on the transition from the kernel regime to the rich regime (Lyu et al., 2024; Kumar et al., 2024). Our study provides a novel perspective based on the emergence of neural collapse, offering a new connection to the representation-learning view of Liu et al. (2022).
@@
The IB principle (Tishby et al., 2000) has been analyzed in the context of deep learning by modeling the successive neural network layers as a Markov chain (Tishby & Zaslavsky, 2015; Shwartz-Ziv & Tishby, 2017; Saxe et al., 2018; Geiger, 2021; Lorenzen et al., 2022; Adilova et al., 2023; Butakov et al., 2024a, b). In addition to these studies analyzing DNN training from the IB perspective, minimizing the IB objective has also been shown to benefit generalization bounds (Kawaguchi et al., 2023; Sefidgaran et al., 2023). Several studies have attempted to explain the IB dynamics. For example, Shwartz-Ziv et al. (2019) attributed the compression phase to the diffusion component of stochastic gradient descent (SGD), while Goldfeld et al. (2019) and Koch & Ghosh (2025) discussed its relation to geometric compression and grokking, respectively. However, these studies lack the rigorous theoretical analysis of the IB dynamics, while our work offers a new explanation based on neural collapse, a geometric form of compression.
@@
Neural collapse (Papyan et al., 2020) is a late-stage training phenomenon where the representations of the training data converge to a simplex equiangular tight frame (ETF) formed with the class mean representations. A major line of research analyzes the optimality of such solutions under an unconstrained feature model (UFM) and its variants (Fang et al., 2021; Mixon et al., 2022; Lu & Steinerberger, 2022; Tirer & Bruna, 2022; Thrampoulidis et al., 2022; Dang et al., 2023; Tirer et al., 2023; Súkeník et al., 2023, 2024; Jiang et al., 2024), where the features are treated directly as optimization variables. Several works have examined the learning dynamics toward neural collapse under UFM (Zhu et al., 2021; Mixon et al., 2022; Ji et al., 2022; Zhou et al., 2022a, b), but this setting cannot reflect the input data properties and diverges from the actual parameter-based training dynamics. Beyond UFM setting, recent studies instead analyze parameter updates under weight regularization, opening a promising new direction (Jacot et al., 2025; Wu & Mondelli, 2025). Building on this, our study broadens the understanding of training dynamics by revealing the connection between late-phase phenomena and neural collapse.
@@ 3 Emergent behavior in late-phase training
We now introduce the population within-class variance, which will play a central role in this section.
@@ 3.1 Grokking: Phase 2 Generalization Dynamics
In the context of grokking, Liu et al. (2022) empirically shows that the acquisition of structured representations leads to a transition from memorization to generalization, which we further analyze here. As a simple observation, since the network output is given by $f({\bm{x}})={\bm{W}}g({\bm{x}})$, achieving good generalization is facilitated when the representations of each class are linearly separable in distribution. When the training loss is sufficiently reduced at some time step $\tau_{1}$, the existence of a weight matrix ${\bm{W}}(\tau_{1})$ indicates that the representations of the training set can be linearly separable; however, this does not guarantee that the representations in distribution are linearly separable. Intuitively, when the representations of each class are more tightly clustered, achieving linear separability becomes easier, and better generalization performance can be expected. Formalizing this intuition yields the following theorem.
@@
This theorem follows directly from applying the union bound and a variance-based tail probability bound. The proof is given in Section A.1. Theorem 3.2 states that improving test accuracy is facilitated by two conditions: i) the class mean representations $\mathbb{E}_{X|Y=c}\left[\tilde{g}(X)\right]$ become more aligned with the corresponding last-layer weights ${\bm{w}}_{c}$, and ii) the population within-class variance decreases. From a grokking perspective, when the training loss is sufficiently reduced, it remains unclear how well the classifier $\bm{W}$} aligns with the class means of the representations. Nevertheless, independent of the properties of $\bm{W}$}, reducing the within-class variance of the representations tightens the upper bound on the generalization error. To uncover the mechanism of grokking, it remains to show that the within-class variance decreases even after the training accuracy has been sufficiently improved.
@@ 3.2 IB Dynamics
where $\beta>0$ is a trade-off parameter controlling the balance between compression and information preservation. When analyzing DNNs within the IB framework, we encounter several difficulties: for a deterministic network and a continuous representation, $I(Z;X)$ can be infinite, making the analysis ill-defined; furthermore, the network parameters are not reflected in the IB analysis (Saxe et al., 2018; Amjad & Geiger, 2019; Goldfeld et al., 2019). To address these issues, we analyze the representation after independently adding an arbitrarily small Gaussian noise $E\sim N(\bm{0},\sigma^{2}{\bm{I}})$, $\sigma\ll 1$, which is a standard approach in IB analysis of DNNs (Saxe et al., 2018; Goldfeld et al., 2019; Butakov et al., 2024b). It should be noted that such a small amount of noise has a negligible effect on the network’s output, making it a good proxy for the actual network. Since we are currently focusing on the output of the feature extractor $g$, we denote the representation as $Z=g(X)+B_{g}E$, where $B_{g}=\sup_{{\bm{x}}\in{\mathcal{X}}}\|g({\bm{x}})\|_{2}$ denotes the output scale of $g$ introduced in Definition 3.1. We use $\bm{z}$} to denote the realization of $Z$.
Representation dynamics in the two-dimensional $(I(Z;X),I(Z;Y))$ plane, which is called information plane, is a useful tool for analyzing the training dynamics of DNNs from the IB perspective (Shwartz-Ziv & Tishby, 2017). As a preliminary observation, to perform classification accurately with a neural network, both $I(Z;X)$ and $I(Z;Y)$ need to be sufficiently large; this follows from the following proposition (see Section A.2 for the proof):
@@
The proof is given in Section A.2. Theorem 3.4 highlights the role of within-class variance in reducing superfluous information, corresponding to the evolution toward the upper-left in the information plane. Together with Theorem 3.2, the analysis in this section establishes population within-class variance as a key measure. In the next section, we examine how this measure decreases during training and, for grokking, how it proceeds on a different time scale from fitting training set.
@@ 4.1 Population Within-class Variance and Neural Collapse
We first provide a formal bound on the difference between the population within-class variance, which served as an important measure in the previous section, and its empirical counterpart. The following result is based on a standard approach of uniform convergence commonly used in generalization error analysis. The proof is provided in Section A.3.
@@ 4.2 Neural Collapse Dynamics
Although we do not analyze NC2 directly, we define it as the condition number of the matrix of class mean vectors, $\mathrm{NC2}=\kappa\left(\begin{pmatrix}{\bm{\mu}}_{1},\ldots,{\bm{\mu}}_{K}\end{pmatrix}\right)$ and examine its behavior in the experimental section.
@@
We now analyze the dynamics of neural collapse measured in terms of RNC1, together with the convergence of the training loss. Our results build on the recent results of Jacot et al. (2025), which assume (i) a pyramidal network architecture, (ii) a smooth activation function, and (iii) a specific initialization condition; their formal statements are given in Section A.4. The next theorem establishes that the training loss and RNC1 converge on different time scales. Here we use the standard Big-O notations $O(\cdot)$ and $\Omega(\cdot)$ to describe the dependence on $\lambda$, $\eta$, $\epsilon_{1}$, and $\epsilon_{2}$.
@@
Suppose that the network $f$ satisfies Assumptions A.3, A.4 and A.5 and that the input domain $\mathcal{X}$} is bounded. Fix $0<\epsilon_{1}<\frac{1}{8}\min_{c\in[K]}n_{c}$ and $\epsilon_{2}>0$. For weight decay $\lambda=O(\epsilon_{1})$, learning rate $\eta=O(\epsilon_{2})$, and time steps $\tau_{1}<\tau_{2}$ satisfying
@@ 5.1 Grokking
We first analyze the relationship between grokking, within-class variance, and neural collapse. Following Liu et al. (2023a), we train an MLP on the MNIST dataset (LeCun et al., 2010). The model has four layers with architecture $[784,200,200,200,10]$ and ReLU activation. The initialization scale is increased by a factor of $8$ as in Liu et al. (2023a), and we use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $1\mathrm{e}{-3}$ and weight decay of $0.01$. Additional results on other datasets and architectures, including convolutional neural networks (CNNs) and Transformers (Vaswani et al., 2017), are presented in Section C.1.
@@
Figure 2: Margins of individual examples at two time steps during grokking. The margin of each example is defined as the signed distance from its representation to the decision boundary determined by the last-layer classifier, calculated as $\left(\langle{\bm{w}}_{0}-{\bm{w}}_{1},g({\bm{x}})\rangle+b_{0}-b_{1}\right)/\|{\bm{w}}_{0}-{\bm{w}}_{1}\|_{2}$, where $b_{c}$ denotes the bias term for class $c$. We trained a 4-layer MLP on the MNIST dataset. These results reveal the link between representation variance and generalization, supporting Theorems 3.2 and 4.1.

### v2

@@ 1 Introduction
Deep neural networks (DNNs) have demonstrated remarkable success across a variety of tasks, including computer vision, natural language processing, and reinforcement learning, yet their training dynamics under gradient descent often reveal unexpected behavior. In particular, when training continues beyond the point where the training loss has been sufficiently reduced, several intriguing late-phase phenomena have been reported. One such phenomenon is grokking (Power et al., 2022), where models initially converge to an overfitting solution that perfectly fits the training data but fails to generalize to unseen data. However, when training continues for a sufficiently long time, the models unexpectedly generalize. Another example is the information bottleneck (IB) framework (Tishby et al., 2000; Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017), an information-theoretic perspective on representation learning in DNNs that formalizes the goal of retaining task-relevant information while compressing the task-irrelevant input information. One particularly intriguing observation here is that DNNs do not move directly toward an IB-optimal solution; instead, they first enter a fitting phase where they memorize the training data, followed by a later compression phase in the later training stage during which task-irrelevant input information is discarded.
@@
We consider a $K$-class classification problem with a dataset $S=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ consisting of $N$ examples, where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ is the input and $y_{i}\in[K](\coloneqq\{1,2,\dots,K\})$ is the class label. The input domain is ${\mathcal{X}}\subseteq\mathbb{R}^{d}$, and ${\bm{y}}_{i}$ denotes the one-hot encoding label in $\{0,1\}^{K}$. Let ${\bm{X}}\in\mathbb{R}^{d\times N}$ and ${\bm{Y}}\in\{0,1\}^{K\times N}$ denote the training inputs and one-hot labels, respectively. We use $S_{c}=\{({\bm{x}}_{i},y_{i})\in S\mid y_{i}=c\}$ to denote the subset of class-$c$ examples, and, by abuse of notation, also the corresponding index set $\{i\in[N]\mid y_{i}=c\}$. We denote its cardinality as $n_{c}=|S_{c}|$. Let $X$ and $Y$ denote the random variables representing the input and label, respectively. We write $p_{X,Y}$ for the joint distribution over $(X,Y)$, and $p_{X}$ and $p_{Y}$ for their marginal distributions. When the variables are clear from the context, we simply write $p$ to refer to the corresponding distribution. Let $I(X;Y)$ be the mutual information (MI) between random variables $X$ and $Y$. We denote the multivariate Gaussian distribution with mean $\bm{\mu}$} and covariance matrix $\bm{\Sigma}$} as $N({\bm{\mu}},{\bm{\Sigma}})$. We use $\log(\cdot)$ to denote the natural logarithm. For linear-algebraic notation, for a matrix $\bm{A}$}, we use $\|{\bm{A}}\|_{2}$ to denote the spectral norm, and $\|{\bm{A}}\|_{2,1}$ to denote the $(2,1)$ matrix norm, defined as $\|{\bm{A}}\|_{2,1}=\sum_{i}\|{\bm{a}}_{i}\|_{2}$, where ${\bm{a}}_{i}$ is the $i$-th column of $\bm{A}$}. Finally, we use $\langle{\bm{a}},{\bm{b}}\rangle={\bm{a}}^{\top}{\bm{b}}$ to denote the standard Euclidean inner product.
@@ 2 Related Work
On the empirical side, several studies attempted to explain the cause of grokking in terms of the parameter compression (Liu et al., 2023a; Varma et al., 2023) and some complexity measures (Nanda et al., 2023; Liu et al., 2023b; Humayun et al., 2024; DeMoss et al., 2025). Other studies sought to relate grokking to other training-related concepts, such as double descent (Davies et al., 2023; Huang et al., 2024) and optimization stability (Thilak et al., 2022). On the theoretical side, a high-dimensional limit of the linear model was analyzed in the setting of regression (Levi et al., 2024) and binary classification (Beck et al., 2025). There are also studies analyzing two-layer networks with XOR data (Xu et al., 2024) and mean-field analysis (Rubin et al., 2024). The most closely related line of work focuses on the transition from the kernel regime to the rich regime (Lyu et al., 2024; Kumar et al., 2024). Our study provides a novel perspective based on the emergence of neural collapse, offering a new connection to the representation-learning view of Liu et al. (2022b).
@@
The IB principle (Tishby et al., 2000) has been analyzed in the context of deep learning by modeling the successive neural network layers as a Markov chain (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017; Saxe et al., 2018; Geiger, 2021; Lorenzen et al., 2022; Adilova et al., 2023; Butakov et al., 2024a; b). In addition to these studies analyzing DNN training from the IB perspective, minimizing the IB objective has also been shown to benefit generalization bounds (Kawaguchi et al., 2023; Sefidgaran et al., 2023). Several studies have attempted to explain the IB dynamics. For example, Shwartz-Ziv et al. (2019) attributed the compression phase to the diffusion component of stochastic gradient descent (SGD), while Goldfeld et al. (2019) and Koch and Ghosh (2025) discussed its relation to geometric compression and grokking, respectively. However, these studies lack the rigorous theoretical analysis of the IB dynamics. A related line of work analyzes late-phase behavior through reconstruction loss rather than mutual information (Schneider and Prabhushankar, 2024; Schneider, 2025), though this perspective provides only an indirect proxy for the IB dynamics. In contrast, our work offers a new explanation of IB dynamics based on neural collapse, a geometric form of compression.
@@
Neural collapse (Papyan et al., 2020) is a late-stage training phenomenon where the representations of the training data converge to a simplex equiangular tight frame (ETF) formed with the class mean representations. A major line of research analyzes the optimality of such solutions under an unconstrained feature model (UFM) and its variants (Fang et al., 2021; Mixon et al., 2022; Lu and Steinerberger, 2022; Tirer and Bruna, 2022; Thrampoulidis et al., 2022; Dang et al., 2023; Tirer et al., 2023; Súkeník et al., 2023; 2024; Jiang et al., 2024), where the features are treated directly as optimization variables. Several works have examined the learning dynamics toward neural collapse under UFM (Zhu et al., 2021; Mixon et al., 2022; Ji et al., 2022; Zhou et al., 2022a; b), but this setting cannot reflect the input data properties and diverges from the actual parameter-based training dynamics. Beyond UFM setting, recent studies instead analyze parameter updates under weight regularization, opening a promising new direction (Jacot et al., 2025; Wu and Mondelli, 2025). Building on this, our study broadens the understanding of training dynamics by revealing the connection between late-phase phenomena and neural collapse.
@@ 3 Emergent behavior in late-phase training
We now introduce the population within-class variance, which will play a central role in this section. We remark that this metric is not introduced as an ad-hoc measure; rather, it naturally arises from our analysis of the second-phase dynamics presented below. In particular, the scale-invariant formulation in Definition 3.1 ensures that this quantity reflects genuine geometric concentration, disentangled from changes in output scale. This motivates the following definition.
@@ 3.1 Grokking: Phase 2 Generalization Dynamics
In the context of grokking, Liu et al. (2022b) empirically shows that the acquisition of structured representations leads to a transition from memorization to generalization, which we further analyze here. As a simple observation, since the network output is given by $f({\bm{x}})={\bm{W}}g({\bm{x}})$, achieving good generalization is facilitated when the representations of each class are linearly separable in distribution. When the training loss is sufficiently reduced at some time step $\tau_{1}$, the existence of a weight matrix ${\bm{W}}(\tau_{1})$ indicates that the representations of the training set can be linearly separable; however, this does not guarantee that the representations in distribution are linearly separable. Intuitively, when the representations of each class are more tightly clustered, achieving linear separability becomes easier, and better generalization performance can be expected. Formalizing this intuition yields the following theorem.
@@
This theorem follows directly from applying the union bound and a variance-based tail probability bound. The proof is given in Section B.1. Theorem 3.2 states that improving test accuracy is facilitated by two conditions: i) the class mean representations $\mathbb{E}_{X|Y=c}\left[\tilde{g}(X)\right]$ become more aligned with the corresponding last-layer weights ${\bm{w}}_{c}$, and ii) the population within-class variance decreases. From a grokking perspective, when the training loss is sufficiently reduced, it remains unclear how well the classifier $\bm{W}$} aligns with the class means of the representations. Nevertheless, independent of the properties of $\bm{W}$}, reducing the within-class variance of the representations tightens the upper bound on the generalization error. To uncover the mechanism of grokking, it remains to show that the within-class variance decreases even after the training accuracy has been sufficiently improved.
@@ 3.2 IB Dynamics
where $\beta>0$ is a trade-off parameter controlling the balance between compression and information preservation. When analyzing DNNs within the IB framework, we encounter several difficulties: for a deterministic network and a continuous representation, $I(Z;X)$ can be infinite, making the analysis ill-defined; furthermore, the network parameters are not reflected in the IB analysis (Saxe et al., 2018; Amjad and Geiger, 2019; Goldfeld et al., 2019). To address these issues, we analyze the representation after independently adding an arbitrarily small Gaussian noise $E\sim N(\bm{0},\sigma^{2}{\bm{I}})$, $\sigma\ll 1$, which is a standard approach in IB analysis of DNNs (Saxe et al., 2018; Goldfeld et al., 2019; Butakov et al., 2024b). It should be noted that such a small amount of noise has a negligible effect on the network’s output, making it a good proxy for the actual network. Since we are currently focusing on the output of the feature extractor $g$, we denote the representation as $Z=g(X)+B_{g}E$, where $B_{g}=\sup_{{\bm{x}}\in{\mathcal{X}}}\|g({\bm{x}})\|_{2}$ denotes the output scale of $g$ introduced in Definition 3.1. We use $\bm{z}$} to denote the realization of $Z$.
Representation dynamics in the two-dimensional $(I(Z;X),I(Z;Y))$ plane, which is called information plane, is a useful tool for analyzing the training dynamics of DNNs from the IB perspective (Shwartz-Ziv and Tishby, 2017). As a preliminary observation, to perform classification accurately with a neural network, both $I(Z;X)$ and $I(Z;Y)$ need to be sufficiently large; this follows from the following proposition (see Section B.2 for the proof):
@@
The proof is given in Section B.2. We further show the tightness and behavior of this upper bound in Proposition B.3. Theorem 3.4 highlights the role of within-class variance in reducing superfluous information, corresponding to the evolution toward the upper-left in the information plane. Together with Theorem 3.2, the analysis in this section establishes population within-class variance as a key measure. In the next section, we examine how this measure decreases during training and, for grokking, how it proceeds on a different time scale from fitting training set.
@@ 4.1 Population Within-class Variance and Neural Collapse
We first provide a formal bound on the difference between the population within-class variance, which served as an important measure in the previous section, and its empirical counterpart. The following result is based on a standard approach of uniform convergence commonly used in generalization error analysis. The proof is provided in Section B.3.
@@ 4.2 Neural Collapse Dynamics
We now analyze the dynamics of neural collapse measured in terms of RNC1, together with the convergence of the training loss. Our results build on the recent results of Jacot et al. (2025), which assume (i) a pyramidal network architecture, (ii) a smooth activation function, and (iii) a specific initialization condition; their formal statements are given in Section B.4. The next theorem establishes that the training loss and RNC1 converge on different time scales. Here we use the standard Big-O notations $O(\cdot)$ and $\Omega(\cdot)$ to describe the dependence on $\lambda$, $\eta$, $\epsilon_{1}$, and $\epsilon_{2}$.
@@
Suppose that the network $f$ satisfies Assumptions B.3, B.4 and B.5 and that the input domain $\mathcal{X}$} is bounded. Fix $0<\epsilon_{1}<\frac{1}{8}\min_{c\in[K]}n_{c}$ and $\epsilon_{2}>0$. For weight decay $\lambda=O(\epsilon_{1})$, learning rate $\eta=O(\epsilon_{2})$, and time steps $\tau_{1}<\tau_{2}$ satisfying
@@ 5.1 Grokking
We first analyze the relationship between grokking, within-class variance, and neural collapse. Following Liu et al. (2023a), we train an MLP on the MNIST dataset (LeCun et al., 2010). The model has four layers with architecture $[784,200,200,200,10]$ and ReLU activation. The initialization scale is increased by a factor of $8$ as in Liu et al. (2023a), and we use the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $1\mathrm{e}{-3}$ and weight decay of $0.01$. Additional results on other datasets and architectures, including convolutional neural networks (CNNs) and Transformers (Vaswani et al., 2017), are presented in Section D.1. In examining the dynamics of neural collapse, we primarily track the RNC1 score, which is the focus of our theoretical analysis. As an additional geometric metric, we also report the NC2 score, another geometric aspect of neural collapse. We define it as the condition number of the matrix of class mean vectors, $\mathrm{NC2}=\kappa\left(\begin{pmatrix}{\bm{\mu}}_{1},\ldots,{\bm{\mu}}_{K}\end{pmatrix}\right)$. We include NC2 to provide supplementary geometric insight into the arrangement of the class mean vectors.
@@
Figure 2: Margins of individual examples at two time steps during grokking. The margin of each example is defined as the signed distance from its representation to the decision boundary determined by the last-layer classifier, calculated as $\left(\langle{\bm{w}}_{0}-{\bm{w}}_{1},g({\bm{x}})\rangle+b_{0}-b_{1}\right)/\|{\bm{w}}_{0}-{\bm{w}}_{1}\|_{2}$, where $b_{c}$ denotes the bias term for class $c$. We trained a 4-layer MLP on the MNIST dataset. These results reveal the link between representation variance and generalization, supporting Theorems 3.2 and 4.1. We additionally provide similar visualizations for several other class pairs in Section D.1.